A newly issued bond has a maturity of 10 years and pays a 5.4% coupon rate (with coupon payments coming once annually). The bond sells at par value. a. What are the convexity and the duration of the bond? Use the formula for convexity in footnote 7. (Round your answers to 3 decimal places.) b. Find the actual price of the bond assuming that its yield to maturity immediately increases from 5.4% to 6.4% (with maturity still 10 years). Assume a par value of 100. (Round your answer to 2 decimal places.) c. What price would be predicted by the modified duration rule ΔPP=−D*Δy? What is the percentage error of that rule? (Negative answers should be indicated by a minus sign. Round your answers to 2 decimal places.) d. What price would be predicted by the modified duration-with-convexity rule ΔPP=−D*Δy+12×Convexity×(Δy)2? What is the percentage error of that rule? (Negative answers should be indicated by a minus sign. Round your answers to 2 decimal places.)
Answer #1

a

phpOUK6FH.png

C =

Period Cash Flow Discounting factor PV Cash Flow Duration Calc Convexity Calc
0 ($100.00) =(1+YTM/number of coupon payments in the year)^period =cashflow/discounting factor =PV cashflow*period =duration calc*(1+period)/(1+YTM/N)^2
1                   5.40                                                             1.05                      5.12                    5.12                    9.22
2                   5.40                                                             1.11                      4.86                    9.72                  26.25
3                   5.40                                                             1.17                      4.61                  13.84                  49.82
4                   5.40                                                             1.23                      4.38                  17.50                  78.77
5                   5.40                                                             1.30                      4.15                  20.76                112.11
6                   5.40                                                             1.37                      3.94                  23.63                148.91
7                   5.40                                                             1.45                      3.74                  26.16                188.37
8                   5.40                                                             1.52                      3.55                  28.36                229.78
9                   5.40                                                             1.61                      3.36                  30.27                272.51
10              105.40                                                             1.69                    62.29                622.92              6,168.02
      Total                798.29              7,283.77
Macaulay duration =(∑ Duration calc)/(bond price*number of coupon per year)
=798.29/(100*1)
=7.982904
Modified duration = Macaulay duration/(1+YTM)
=7.98/(1+0.054)
=7.574
Convexity =(∑ convexity calc)/(bond price*number of coupon per year^2)
=7283.77/(100*1^2)
=72.838

b

Actual bond price change
                  K = N
Bond Price =∑ [( Coupon)/(1 + YTM)^k]     +   Par value/(1 + YTM)^N
                   k=1
                  K =10
Bond Price =∑ [(5.4*100/100)/(1 + 6.4/100)^k]     +   100/(1 + 6.4/100)^10
                   k=1
Bond Price = 92.78

c

Using only modified duration
Mod.duration prediction = -Mod. Duration*Yield_Change*Bond_Price
=-7.57*0.01*100
=-7.57
%age change in bond price=Mod.duration prediction/bond price
=-7.57/100
=-7.57%
New bond price = bond price+Modified duration prediction
=100-7.57
=92.43
Difference in price predicted and actual
=predicted price-actual price
=92.43-92.78
=-0.35
%age difference = difference/actual-1
=-0.35/92.78
=-0.38%

d

Using convexity adjustment to modified duration
Convexity adjustment = 0.5*convexity*Yield_Change^2*Bond_Price
0.5*72.84*0.01^2*100
=0.36
%age change in bond price=(Mod.duration pred.+convex. Adj.)/bond price
=(-7.57+0.36)/100
=-7.21%
New bond price = bond price+Mod.duration pred.+convex. Adj.
=100-7.57+0.36
=92.79
Difference in price predicted and actual
=predicted price-actual price
=92.79-92.78
=0.01
%age difference = difference/actual-1
=0.01/92.78
=0.01%